# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v10

@@ 5.3 Theorem on TVD Upper Bounds
Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have (Kullback and Leibler, 1951):
@@
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance. By Pinsker’s inequality (Pinsker, 1964), we have: …
@@
Theorem 4 (TVD Upper Bound for Arbitrary Prediction Strategies with Noisy Inputs and Uncovered Training Regions) Let the base space $X$ be a finite set. The training information domain of the model $M$ is $S_{ou}\subseteq X$. The true problem domain is $S_{oq}\subseteq X$, and the true distribution $p_{q}$ satisfies $\sum_{x\in S_{oq}}{p_{q}(x)}=1$. The noisy information ontological state $\SoqSoq+\widetilde{S_{oq}}=S_{oq}\backslash$ $\cup S_{oq}^{+}$, as defined in Definition 3, replaces $S_{oq}$ as the actual input to the model $M$. The noise loss probability mass is $P_{l}=\sum_{x\in S_{oq}^{-}}{p_{q}(x)}\in[0,1]$, and the noise addition probability mass is $P_{a}=\sum_{x\in S_{oq}^{+}}{p^{+}(x)}\in[0,1]$, where $p^{+}$ is a non-negative kernel on the noise addition region, satisfying $1-P_{l}+P_{a}>0$. The distribution of the noisy input $\widetilde{S_{oq}}$ is given by $\widetilde{p_{q}}(x)=(p_{q}(x)\bullet 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\bullet 1\left\{x\in S_{oq}^{+}\right\})/(1-P_{l}+P_{a})$. Let $V=\widetilde{S_{oq}}\backslash S_{ou}$ denote the region not covered by training information, with uncovered probability mass $P_{uncov}=\sum_{x\in V}{\widetilde{p_{q}}(x)}\in[0,1]$. The conditional distribution on $V$ is $r(x)=\widetilde{p_{q}}(x)/P_{uncov}$ (and when $P_{uncov}=0$, all terms involving $V$ below are taken as 0). The predictive strategy $p_{0}$ on the uncovered region $V$ is an arbitrary distribution, satisfying $\sum_{x\in V}{p_{0}(x)=1,}p_{0}(x)\geq 0$. The model distribution $p_{M}$ is determined by training on the covered region $U=\widetilde{S_{oq}}\cap S_{ou}$, and on $V$, $p_{M}(x)=P_{uncov}p_{0}(x)$. Assume that the model's fitting error on the covered region is $E_{fit}=D_{KL}\left(\widetilde{p_{q}}|_{U}\parallel p_{M}|_{U}\right)$, where $\widetilde{p_{q}}|_{U}$ and $p_{M}|_{U}$ denote the distributions restricted to $U$ and normalized (under the ideal conditions assumed in Theorem 3, $E_{fit}$ tends to zero).
@@
$\widetilde{p_{q}}(x)=\frac{p_{q}(x)\bullet 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\bullet 1\left\{x\in S_{oq}^{+}\right\}}{1-P_{l}+P_{a}}$ (5.19)
@@ 6.4.3 Partition-Based Explanation
Fig. 6(a) presents an overall view of numerous inputs processed and predicted by model M. This visualization also represents an ontological state of M as it handles a wide range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping. Thus, according to Definition 4, Figure 5(a) constitutes a valid explanation for the distribution of prediction polarities by M.
@@ 7 Discussion and Conclusion
Appendix A Appendix A. Formal Representation of States
@@
Appendix B Appendix B. Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata (Xu, 2024).
@@
Appendix C Appendix C. Model Instance Validation of MLT-MF
@@ C.1 A Simple Feedforward Neural Network Model
We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Fig. C.1). Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
Figure 7: Fig. C.1 Schematic diagram of feedforward neural network model M
@@ C.2 Symbol Definitions
MLT-MF uses a formal system to express the operational logic of each phase throughout the model’s full lifecycle. Defining the symbols is a prerequisite for formal modeling. To this end, Tables C.1–C.4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
@@ C.3 Formal Representations of Information at Each Stage of Model M
At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Fig. C.1. Although M is a simple model, it contains all the essential elements of a neural network, and the discussion has not relied on any special-case treatment of M. This demonstrates that MLT-MF is capable of supporting the construction of well-formed, formal representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.

### v11

@@ 5.3 Theorem on TVD Upper Bounds
Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have (Goodfellow et al., 2016):
@@
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance. By Pinsker’s inequality (Pinsker, 1964), we have:
@@
Theorem 4 (TVD Upper Bound for Arbitrary Prediction Strategies with Noisy Inputs and Uncovered Training Regions) Let the base space $X$ be a finite set. The training information domain of the model $M$ is $S_{ou}\subseteq X$. The true problem domain is $S_{oq}\subseteq X$, and the true distribution $p_{q}$ satisfies $\sum_{x\in S_{oq}}{p_{q}(x)}=1$. The noisy information ontological state $\SoqSoq+\widetilde{S_{oq}}=S_{oq}\backslash$ $\cup S_{oq}^{+}$, as defined in Definition 3, replaces $S_{oq}$ as the actual input to the model $M$. The noise loss probability mass is $P_{l}=\sum_{x\in S_{oq}^{-}}{p_{q}(x)}\in[0,1]$, and the noise addition probability mass is $P_{a}=\sum_{x\in S_{oq}^{+}}{p^{+}(x)}\in[0,1]$, where $p^{+}$ is a non-negative kernel on the noise addition region, satisfying $1-P_{l}+P_{a}>0$. The distribution of the noisy input $\widetilde{S_{oq}}$ is given by $\widetilde{p_{q}}(x)=(p_{q}(x)\cdot 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\cdot 1\left\{x\in S_{oq}^{+}\right\})/(1-P_{l}+P_{a})$. Let $V=\widetilde{S_{oq}}\backslash S_{ou}$ denote the region not covered by training information, with uncovered probability mass $P_{uncov}=\sum_{x\in V}{\widetilde{p_{q}}(x)}\in[0,1]$. The conditional distribution on $V$ is $r(x)=\widetilde{p_{q}}(x)/P_{uncov}$ (and when $P_{uncov}=0$, all terms involving $V$ below are taken as 0). The predictive strategy $p_{0}$ on the uncovered region $V$ is an arbitrary distribution, satisfying $\sum_{x\in V}{p_{0}(x)=1,}p_{0}(x)\geq 0$. The model distribution $p_{M}$ is determined by training on the covered region $U=\widetilde{S_{oq}}\cap S_{ou}$, and on $V$, $p_{M}(x)=P_{uncov}p_{0}(x)$. Assume that the model's fitting error on the covered region is $E_{fit}=D_{KL}\left(\widetilde{p_{q}}|_{U}\parallel p_{M}|_{U}\right)$, where $\widetilde{p_{q}}|_{U}$ and $p_{M}|_{U}$ denote the distributions restricted to $U$ and normalized (under the ideal conditions assumed in Theorem 3, $E_{fit}$ tends to zero).
@@
$\widetilde{p_{q}}(x)=\frac{p_{q}(x)\cdot 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\cdot 1\left\{x\in S_{oq}^{+}\right\}}{1-P_{l}+P_{a}}$ (5.19)
@@ 6.4.3 Partition-Based Explanation
Fig. 6(a) presents an overall view of numerous inputs processed and predicted by model M. This visualization also represents an ontological state of M as it handles a wide range of input information—the real processing and prediction procedure of M is a direct result of this ontological mapping. Thus, according to Definition 4, Fig. 6(a) constitutes a valid explanation for the distribution of prediction polarities by M.
@@ 7 Discussion and Conclusion
Appendix A Formal Representation of States
@@
Appendix B Finite Automaton States
Using the formal expression of states, we can study in depth the special and important object of finite automata (Editorial Committee of Dictionary of Mathematics, 2002).
@@ Appendix B Finite Automaton States
Appendix C Model Instance Validation of MLT-MF
@@ C.1 A Simple Feedforward Neural Network Model
We choose a feedforward neural network M with an input layer containing two inputs, a hidden layer containing three neurons, and an output layer containing two outputs (Fig. 7). Its simplicity helps explain the essence of MLT-MF in a clear and straightforward manner.
Figure 7: Schematic diagram of feedforward neural network model M
@@ C.2 Symbol Definitions
MLT-MF uses a formal system to express the operational logic of each phase throughout the model’s full lifecycle. Defining the symbols is a prerequisite for formal modeling. To this end, Tables 2–7 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
@@ C.3 Formal Representations of Information at Each Stage of Model M
At this point, based on MLT-MF, we have constructed the formal well-formed formula representations for all ontological and carrier states of each component of the feedforward neural network model M, as illustrated in Fig. 7. Although M is a simple model, it contains all the essential elements of a neural network, and the discussion has not relied on any special-case treatment of M. This demonstrates that MLT-MF is capable of supporting the construction of well-formed, formal representations for a wide variety of neural networks, providing a solid foundation for detailed analysis and optimization.